OpenStudy (anonymous):

What is the solution to the system of equations? y = –5x + 3 y = 1 (0.4, 1) (0.8, 1) (1, 0.4) (1, 0.8)

OpenStudy (neonumbrella5115):

To solve this system of equations, first you will need to substitute y=1 into the first equation for y.

OpenStudy (anonymous):

I don't understand

jhonyy9 (jhonyy9):

so because there given y=1 again you need calculi just the value of x y=-5x+3 y=1 so this mean that you can writing 1=-5x+3 5x=2 x= 2/5 = 0,4 y=1 x=0,4 y=1 (0,4 ,1) so this mean that first choices is right sure